# Mapping the Cloud-Driven Atmospheric Dynamics & Chemistry of an Isolated Exoplanet Analog with Harmonic Signatures

### v1

@@ Abstract
Young planetary-mass objects and brown dwarfs near the L/T spectral transition exhibit enhanced spectrophotometric variability over field brown dwarfs. Patchy clouds, auroral processes, stratospheric hot spots, and complex carbon chemistry have all been proposed as potential sources of this variability. Using time-resolved, low-to-mid-resolution spectroscopy collected with the JWST/NIRISS and NIRSpec instruments, we apply harmonic analysis to SIMP J013656.5+093347, a highly variable, young, isolated planetary-mass object. Employing Bayesian inference via dynamic nested sampling, we identify fundamental and higher-order harmonics. Odd harmonics () at pressure levels (bar) corresponding to iron (Fe) and forsterite (Mg2SiO4) cloud formation suggest North/South hemispheric asymmetry in the cloudy, and likely equatorial, regions. We use the inferred harmonics, along with 1-D substellar atmospheric models, to map the flux variability by atmospheric pressure level. These vertical maps demonstrate robust interaction between deep convective weather layers and the overlying stratified and radiative atmosphere. We identify distinct time-varying structures in the near-infrared that we interpret as planetary-scale wave (e.g., Rossby or Kelvin)-associated cloud modulation. We detect variability in water (), carbon monoxide (), and methane () molecular signatures. Forsterite cloud modulation is anti-correlated with overlying carbon monoxide and water abundances and correlated with deep methane absorption, suggesting complex interaction between cloud formation, atmospheric chemistry, and temperature structure. Furthermore, we identify distinct harmonic behavior between methane and carbon monoxide absorption bands, providing evidence for time-resolved disequilibrium carbon chemistry. At the lowest pressures (mbar), we find that the mapped methane lines transition from absorption to emission, supporting evidence of high-altitude auroral heating via electron precipitation.
@@ 1 Introduction
Highly variable and rapidly rotating planetary-mass objects and low-mass brown dwarfs provide the opportunity to efficiently observe dynamic gas giant exoplanet-analog atmospheres over multiple rotations, unhindered by host star contamination (Burrows et al., 2001; Faherty et al., 2016). These objects, with inferred masses () overlapping the lowest mass products of stellar formation and the highest mass products of planetary formation, serve as laboratories for understanding how atmospheric circulation, cloud formation, molecular chemistry, and auroral processes interact in H/He-dominated planetary environments.
Brown dwarfs at the transition (K) from mineral cloud-rich late-L dwarfs to the more cloud-sparse mid-T dwarfs exhibit amplified variability (Radigan et al., 2014; Radigan, 2014; Eriksson et al., 2019; Liu et al., 2024). Driving sources of this variability are believed to include multiple, patchy cloud layers perhaps of varying thickness (e.g., Ackerman & Marley, 2001; Burgasser et al., 2002; Reiners & Basri, 2008; Apai et al., 2013); planetary-scale waves (for e.g., Rossby or Kelvin waves, Apai et al., 2017, 2021a; Zhou et al., 2020; Fuda et al., 2024); cloud modulation driven by those planetary-scale waves (Plummer et al., 2024); double-diffusive convection tied to disequilibrium carbon chemistry (Tremblin et al., 2015, 2016); upper atmospheric hot spots (Robinson & Marley, 2014; Morley et al., 2014); dynamically driven temperature variations (Showman & Kaspi, 2013; Nasedkin et al., 2025); and potentially auroral processes Hallinan et al. (2015); Kao et al. (2016, 2018); McCarthy et al. (2025); Nasedkin et al. (2025); Suárez et al. (2025).
@@ 1.1 SIMP J0136
SIMP J013656.5+093347.3 (hereafter SIMP J0136) (Artigau et al., 2006) is an L/T transition object (T) that is highly variable () at near-infrared (NIR) wavelengths (see e.g., Artigau et al., 2009). With a mass ( MJ, Gagné et al., 2017) straddling the minimum deuterium-burning limit, it is often referred to as a planetary-mass object. A relatively young object ( Myr, Gagné et al., 2017) with an inferred surface gravity (Vos et al., 2023), SIMP J0136 fits into a population of younger and lower surface gravity brown dwarfs exhibiting enhanced variability rates when compared to field brown dwarfs (Vos et al., 2022; Liu et al., 2024). SIMP J0136 has a rapid rotation rate ( h, Yang et al., 2016), likely contributing to amplified atmospheric dynamics (e.g., increased wind speeds, Allers et al., 2020), larger equator-to-pole cloud contrasts due to Coriolis effects (Tan & Showman, 2021a, b), and auroral emission due to a strengthened magnetic dynamo and associated currents (e.g., Kao et al., 2018).
@@
Multi-modal observations suggest SIMP J0136’s atmosphere is structured into multiple, time-varying mineral cloud layers. Brewster framework (Burningham et al., 2017) retrievals on spectra spanning m (Burgasser et al., 2008; Sorahana et al., 2013; Filippazzo et al., 2015) infer a thin and patchy forsterite (Mg2SiO4) cloud layer overlying a thicker and more spatially expansive iron (Fe) cloud deck (Vos et al., 2023). Ground-based photometry has detected significant phase shifts between the NIR bands, supporting cloud layers (Yang et al., 2016; McCarthy et al., 2024; Plummer et al., 2024). The suspected forsterite clouds’ phase (inferred via color index reddening) anti-correlation with variation in broadband NIR emission suggest planetary-scale waves may be related to cloud formation and dissipation (Plummer et al., 2024), similar to ammonia/ammonium hydrosulfide cloud modulation in Jupiter’s atmosphere (e.g., Choi et al., 2013; de Pater et al., 2016; Fletcher et al., 2016, 2020).
@@
With its ability to collect spectroscopic and photometric observations across the NIR and mid-IR spectrum, the James Webb Space Telescope (JWST) has revolutionized our understanding of substellar atmospheres like SIMP J0136. Applying a K-means cluster algorithm to time-resolved m spectra resulted in 9 characteristic lightcurve shapes (or clusters) for JWST NIR Spectrograph (NIRSpec, Jakobsen et al., 2022) data and 2 characteristic lightcurves for JWST Mid-IR Instrument (MIRI, Wright et al., 2023) data (McCarthy et al., 2025). These characteristic lightcurves potentially correspond to distinct atmospheric layers (i.e., iron and forsterite cloud layers and cloud-free regions, McCarthy et al., 2025). The K-means algorithm was previously applied to the WISE 1049 AB (L7.5/T0.5) binary brown dwarf system, finding 3 characteristic and spectrophotometrically-resolved lightcurve shapes for the NIRSpec data and 2 characteristic lightcurves for the MIRI data (Biller et al., 2024). A follow-on epoch of observations of WISE 1049 AB with the same instruments found similar results and tied the variation seen in the characteristic lightcurves to patchy clouds and higher-altitude hot spots (Chen et al., 2025).
@@
The JWST Near Infrared Imager and Slitless Spectrograph (NIRISS; Doyon et al. 2023) instrument has also delivered important findings on SIMP J0136. Principal Component Analysis of NIRISS time-series spectroscopy found that rotational variability was driven by spectrally distinct regions (Akhmetshyn et al., 2025), broadly agreeing with McCarthy et al. (2025)’s findings. Multi-wavelength lightcurves collected by NIRISS identified odd harmonics, suggesting asymmetry between Northern and Southern hemispheres, and found initial evidence that broadband emission modulation from forsterite spectral regions was anti-correlated with variability in water and CO absorption lines (Akhmetshyn et al., 2025).
@@
SIMP J0136 possesses a robust magnetic field potentially capable of strong auroral emissions. Radio observations with the Karl G. Jansky Very Large Array at 8-12 GHz detected a G magnetic field (Kao et al., 2018), much stronger than Jupiter’s G equatorial dipole field (see e.g., Barrow, 1960; Acuna & Ness, 1975; Connerney et al., 2018, 2022; Sharan et al., 2022). Auroral heating signatures, in the form of H emission at m, have been detected in NIR observations of Jupiter’s upper atmosphere (see e.g., Kim et al., 1993), but direct detection of H emission remains elusive for T dwarfs like SIMP J0136, and as discussed in Nasedkin et al. (2025), may remain so as H’s quick reaction rate with water may prevent observable H emission (Pineda et al., 2024). Providing further support for Pineda et al. (2024)’s findings, recent observations of the isolated Y dwarf, CWISEP J193518.59-154620.3 (WISE J1935), found emission in the rotational-vibrational methane band (Faherty et al., 2024), suggesting upper atmospheric auroral heating, although H emission was not detected.
@@
Indirect detection methods for inferring the presence of auroral heating are promising. Time-resolved spectral retrievals (via petitRADTRANS, Mollière et al., 2019; Nasedkin et al., 2024) of the same JWST/NIRSpec data presented originally in McCarthy et al. (2025) identify a high-altitude, K temperature inversion, attributed to tentative auroral heating from electron precipitation (Nasedkin et al., 2025).
@@ 1.2 Outline
In this work, we infer and vertically map the atmospheric harmonics in the time-resolved NIR spectra of SIMP J0136. We analyze JWST NIRISS and NIRSpec datasets, collected h apart from one another and originally presented in Akhmetshyn et al. (2025) and McCarthy et al. (2025) respectively, and identify connections between cloud modulation and temperature and chemical structure. The article is structured as follows: §2 describes the observational strategy and data reduction. §3 describes our harmonic model and best-fit scheme. §4 presents the inferred spectral harmonic parameters and uses the best-fit lightcurves and atmospheric models to create vertical (pressure vs. time) heat maps showing the evolution of the atmosphere of SIMP J0136 as well as unique contributions to flux variation from cloud modulation and molecular absorption bands. §5 discusses the implications of our results for the dynamical, temperature, and chemical structure of SIMP J0136 and §6 summarizes our findings.
@@ 2 Observations and Data Reduction
To understand how the atmosphere of SIMP J0136 evolves in the NIR over time, we consider two epochs of observations by two JWST instruments, NIRISS and NIRSpec. NIRISS collections (GTO 1209, PI: Artigau) being made on Jul 22, 2023 from UT 05:08:28 to 09:05:07 (observational cadence: 1.96 min, total observation time: 2.62 h, Akhmetshyn et al., 2025), and NIRSpec observations (GO 3548, PI: Vos) conducted on July 23, 2023 from UT 18:40:56 to 22:05:06 (observational cadence: 1.83 s, total observation time: 2.90 h, McCarthy et al., 2025).
@@
Here we will give a brief overview of the observations and data reduction as they were previously presented in detail in McCarthy et al. (2025) and Akhmetshyn et al. (2025). NIRISS was used in Single Object Slitless Spectroscopy mode (NIRISS/SOSS) to collect 81, spectra from 0.85 to 2.83 m. In this work, we only consider wavelengths m due to background star contamination at shorter wavelengths. The SOSS-Inspired SpectroScopic Extraction (SOSSISSE) pipeline was implemented for data reduction (Lim et al., 2023). NIRSpec BOTS was used in Prism/CLEAR mode to collect 5726, spectra from 0.6 to 5.3 m. As for NIRISS/SOSS above, only wavelengths m are considered. For data reduction, McCarthy et al. (2025) used the JWST Space Telescope Science Institute pipeline version 1.14.0 in default settings.
We use s and s temporal bins for NIRISS/SOSS and NIRSpec/PRISM respectively to create 81 time steps for each data set. It is worth noting that for the NIRSpec/PRISM dataset, this temporal binning reduces the observed % deviation by a factor of versus that seen in McCarthy et al. (2025) and Nasedkin et al. (2025). Both data sets use m spectral binning to compare spectral features at an equivalent resolution. We bin linearly (rather than logarithmically) to more precisely target specific molecular bands, and we acknowledge that the bins will become fractionally smaller at longer wavelengths.
@@ 3.1 Harmonic Model
Using the Python code Imber (Plummer, 2023, 2024) developed in Plummer & Wang (2022, 2023); Plummer et al. (2024), we fit each lightcurve with models of 2nd and 3rd order Fourier modes (or harmonic components), with and without damping. Imber is an open-source code available on GitHub111https://github.com/mkplummer/Imber. The code employs Bayesian methods, specifically dynamic nested sampling (Skilling, 2004, 2006; Higson et al., 2019), to find the best fit for each lightcurve using the Python module Dynesty (Speagle, 2020). Decomposing lightcurves into their harmonic components provides the capability to directly infer parameters including amplitude, period, phase, and damping ratio for each harmonic.
We modify the harmonic model presented in Plummer et al. (2024) by adding a damping term to allow for transient atmospheric dynamics as follows,
@@ 3.2 Fitting Scheme
where denotes inferred parameters, represents observed data, and the applied model. Per Bayesian standard, is the posterior distribution, is the likelihood function, is the prior distribution, and is Bayesian evidence (e.g., Speagle, 2020).
BIC is computed in the traditional manner (Kass & Raftery, 1995),
@@ 3.4 Contribution Function
To map the inferred harmonic modes from wavelength-space to pressure-level space in §4, we use a contribution function. Contribution functions show the emitted radiation at each pressure level in wavelength-space. For this work, we select a cloudless, , Sonora Bobcat (Marley et al., 2021) contribution function (see Figures 3 and 4). The temperature and surface gravity for the model are selected based on findings of Gagné et al. (2017) and Vos et al. (2023). For similar reasons as McCarthy et al. (2025), we make the decision to use a cloudless model to probe the deepest observable pressure levels on SIMP J0136. This approach is supported by the likelihood that planetary-mass objects are composed of an admixture of cloudless and cloudy regions, based on a study of 2MASS 1207 b, an object with similar temperature and surface gravity to SIMP J0136 (Zhang et al., 2025). It should be noted that we do not necessary expect harmonic parameters to increase or decrease in the same manner as the pressure they probe (i.e., the deepest pressure should not necessarily have to have the shortest periods and lowest amplitudes).
As adapting a single contribution function may bias results (e.g., Figure 4 in Dobbs-Dixon & Cowan, 2017), we compared the vertical variation maps described in §4.2.1 to maps created using the retrieved contribution function in Figure 3 of Nasedkin et al. (2025). We found that both maps contained the same key features with only minor differences that did not detract from our conclusions.

### v3

@@ Abstract
Young planetary-mass objects and brown dwarfs near the L/T spectral transition exhibit enhanced spectrophotometric variability over field brown dwarfs. Patchy clouds, auroral processes, stratospheric hot spots, and complex carbon chemistry have all been proposed as potential sources of this variability. Using time-resolved, low-to-mid-resolution spectroscopy collected with the JWST/NIRISS and NIRSpec instruments, we apply harmonic analysis to SIMP J013656.5+093347, a highly variable, young, isolated planetary-mass object. Odd harmonics () at pressure levels (bar) corresponding to iron (Fe) and forsterite (Mg2SiO4) cloud formation suggest potential North/South hemispheric asymmetry in the cloudy, and likely equatorial, regions. We use the inferred harmonics, along with 1-D substellar atmospheric models, to map the flux variability by atmospheric pressure level. These vertical maps demonstrate robust interaction between deep convective weather layers and the overlying stratified and radiative atmosphere. We identify distinct time-varying structures in the near-infrared that we interpret as planetary-scale wave (e.g., Rossby or Kelvin)-associated cloud modulation. We detect deviations from bulk (composite) variability in water (), carbon monoxide (), and methane () molecular signatures. Forsterite cloud modulation is anti-correlated with overlying carbon monoxide and water abundances and correlated with deep methane absorption, suggesting complex interaction between cloud formation, atmospheric chemistry, and temperature structure. Furthermore, we identify distinct harmonic behavior between methane and carbon monoxide absorption bands, providing evidence for time-resolved disequilibrium carbon chemistry. At the lowest pressures (mbar), we find that the mapped methane lines transition from absorption to emission, supporting evidence of high-altitude auroral heating via electron precipitation.
@@ 1 Introduction
Highly variable and rapidly rotating planetary-mass objects and low-mass brown dwarfs provide the opportunity to efficiently observe dynamic gas giant exoplanet-analog atmospheres over multiple rotations, unhindered by host star contamination (Burrows2001; Faherty2016). These objects, with inferred masses () overlapping the lowest mass products of stellar formation and the highest mass products of planetary formation, serve as laboratories for understanding how atmospheric circulation, cloud formation, molecular chemistry, and auroral processes interact in H/He-dominated planetary environments.
Brown dwarfs at the transition (K) from mineral cloud-rich late-L dwarfs to the more cloud-sparse mid-T dwarfs exhibit amplified variability (Radigan2014a; Radigan2014b; Eriksson2019; Liu2024). Driving sources of this variability are believed to include multiple, patchy cloud layers perhaps of varying thickness (e.g., ackerman&marley01; Burgasser2002; Reiners&Basri2008; Apai2013); planetary-scale waves (for e.g., Rossby or Kelvin waves, Apai2017; Apai2021; Zhou2020b; Fuda:2024); cloud modulation driven by those planetary-scale waves (Plummer2024); double-diffusive convection tied to disequilibrium carbon chemistry (Tremblin2015; Tremblin2016); upper atmospheric hot spots (Robinson&Marley2014; Morley2014); dynamically driven temperature variations (showman&kaspi13; Nasedkin2025); and potentially auroral processes Hallinan2015; Kao2016; Kao2018; McCarthy2025; Nasedkin2025; Suarez2025.
@@ 1.1 SIMP J0136
SIMP J013656.5+093347.3 (hereafter SIMP J0136) (Artigau2006) is an L/T transition object (T) that is highly variable () at near-infrared (NIR) wavelengths (see e.g., Artigau2009). With a mass ( MJ, Gagne2017) straddling the minimum deuterium-burning limit, it is often referred to as a planetary-mass object. A relatively young object ( Myr, Gagne2017) with an inferred surface gravity (Vos2023), SIMP J0136 fits into a population of younger and lower surface gravity brown dwarfs exhibiting enhanced variability rates when compared to field brown dwarfs (Vos2022; Liu2024). SIMP J0136 has a rapid rotation rate ( h, Yang2016), likely contributing to amplified atmospheric dynamics (e.g., increased wind speeds, Allers2020), larger equator-to-pole cloud contrasts due to Coriolis effects (tan21a; tan21b), and auroral emission due to a strengthened magnetic dynamo and associated currents (e.g., Kao2018).
@@
Multi-modal observations suggest SIMP J0136’s atmosphere is structured into multiple, time-varying mineral cloud layers. Brewster framework (Burningham2017) retrievals on spectra spanning m (Burgasser2008; Sorahana2013; Filippazzo2015) infer a thin and patchy forsterite (Mg2SiO4) cloud layer overlying a thicker and more spatially expansive iron (Fe) cloud deck (Vos2023). Ground-based photometry has detected phase shifts between the NIR bands, supporting cloud layers (Yang2016; McCarthy:2024; Plummer2024). The suspected forsterite clouds’ phase (inferred via color index reddening) anti-correlation with variation in broadband NIR emission suggest planetary-scale waves may be related to cloud formation and dissipation (Plummer2024), similar to ammonia/ammonium hydrosulfide cloud modulation in Jupiter’s atmosphere (e.g., Choi2013; dePater:2016; Fletcher:2016; Fletcher2020).
@@
With its ability to collect spectroscopic and photometric observations across the NIR and mid-IR spectrum, the James Webb Space Telescope (JWST) has revolutionized our understanding of substellar atmospheres like SIMP J0136. Applying a K-means cluster algorithm to time-resolved m spectra resulted in 9 characteristic lightcurve shapes (or clusters) for JWST NIR Spectrograph (NIRSpec, jakobsen_nirspec_2023) data and 2 characteristic lightcurves for JWST Mid-IR Instrument (MIRI, wright_miri_2023) data (McCarthy2025). These characteristic lightcurves potentially correspond to distinct atmospheric layers (i.e., iron and forsterite cloud layers and cloud-free regions, McCarthy2025). The K-means algorithm was previously applied to the WISE 1049 AB (L7.5/T0.5) binary brown dwarf system, finding 3 characteristic and spectrophotometrically-resolved lightcurve shapes for the NIRSpec data and 2 characteristic lightcurves for the MIRI data (Biller2024). A follow-on epoch of observations of WISE 1049 AB with the same instruments found similar results and tied the variation seen in the characteristic lightcurves to patchy clouds and higher-altitude hot spots (Chen2025).
@@
The JWST Near Infrared Imager and Slitless Spectrograph (NIRISS; doyon_niriss_2023) instrument has also delivered important findings on SIMP J0136. Principal Component Analysis of NIRISS time-series spectroscopy found that rotational variability was driven by spectrally distinct regions (Akhmetshyn2025), broadly agreeing with McCarthy2025’s findings. Multi-wavelength lightcurves collected by NIRISS identified odd harmonics, suggesting asymmetry between Northern and Southern hemispheres, and found initial evidence that broadband emission modulation from forsterite spectral regions was anti-correlated with variability in water and CO absorption lines (Akhmetshyn2025).
@@
SIMP J0136 possesses a robust magnetic field potentially capable of strong auroral emissions. Radio observations with the Karl G. Jansky Very Large Array at 8-12 GHz detected a G magnetic field (Kao2018), much stronger than Jupiter’s G equatorial dipole field (see e.g., Barrow1960; Acuna&Ness1975; Connerney2018; Connerney2022; Sharan2022). Auroral heating signatures, in the form of H emission at m, have been detected in NIR observations of Jupiter’s upper atmosphere (see e.g., Kim1993), but direct detection of H emission remains elusive for T dwarfs like SIMP J0136, and as discussed in Nasedkin2025, may remain so as H’s quick reaction rate with water may prevent observable H emission (Pineda2024). Providing further support for Pineda2024’s findings, recent observations of the isolated Y dwarf, CWISEP J193518.59-154620.3 (WISE J1935), found emission in the rotational-vibrational methane band (Faherty2024), suggesting upper atmospheric auroral heating, although H emission was not detected.
@@
Indirect detection methods for inferring the presence of auroral heating are promising. Time-resolved spectral retrievals (via petitRADTRANS, Molliere2019; Nasedkin2024) of the same JWST/NIRSpec data presented originally in McCarthy2025 identify a high-altitude, K temperature inversion, attributed to tentative auroral heating from electron precipitation (Nasedkin2025).
@@ 1.2 Outline
In this work, we infer and vertically map the atmospheric harmonics in the time-resolved NIR spectra of SIMP J0136. We analyze JWST NIRISS and NIRSpec datasets, collected h apart from one another and originally presented in Akhmetshyn2025 and McCarthy2025 respectively, and identify connections between cloud modulation and temperature and chemical structure. The article is structured as follows: §2 describes the observational strategy and data reduction. §3 describes our harmonic model and best-fit scheme. §4 presents the inferred spectral harmonic parameters and uses the best-fit lightcurves and atmospheric models to create vertical (pressure vs. time) heat maps showing the evolution of the atmosphere of SIMP J0136 as well as spectral features’ deviations from bulk atmospheric variability. §5 discusses the implications of our results for the dynamical, temperature, and chemical structure of SIMP J0136 and §6 summarizes our findings.
@@ 2 Observations and Data Reduction
To understand how the atmosphere of SIMP J0136 evolves in the NIR over time, we consider two epochs of observations by two JWST instruments, NIRISS and NIRSpec. NIRISS collections (GTO 1209, PI: Artigau) being made on Jul 22, 2023 from UT 05:08:28 to 09:05:07 (observational cadence: 1.96 min, total observation time: 2.62 h, Akhmetshyn2025), and NIRSpec observations (GO 3548, PI: Vos) conducted on July 23, 2023 from UT 18:40:56 to 22:05:06 (observational cadence: 1.83 s, total observation time: 2.90 h, McCarthy2025).
@@
Here we will give a brief overview of the observations and data reduction as they were previously presented in detail in McCarthy2025 and Akhmetshyn2025. NIRISS was used in Single Object Slitless Spectroscopy mode (NIRISS/SOSS) to collect 81, spectra from 0.85 to 2.83 m. In this work, we only consider wavelengths m due to background star contamination at shorter wavelengths. The SOSS-Inspired SpectroScopic Extraction (SOSSISSE) pipeline was implemented for data reduction (Lim2023). NIRSpec BOTS was used in Prism/CLEAR mode to collect 5726, spectra from 0.6 to 5.3 m. As for NIRISS/SOSS above, only wavelengths m are considered. For data reduction, McCarthy2025 used the JWST Space Telescope Science Institute pipeline version 1.14.0 in default settings.
We use s and s temporal bins for NIRISS/SOSS and NIRSpec/PRISM respectively to create 81 time steps for each data set. It is worth noting that for the NIRSpec/PRISM dataset, this temporal binning reduces the observed % deviation by a factor of versus that seen in McCarthy2025 and Nasedkin2025. Both data sets use m spectral binning to compare spectral features at an equivalent resolution. We bin linearly (rather than logarithmically) to more precisely target specific molecular bands, and we acknowledge that the bins will become fractionally smaller at longer wavelengths.
@@ 3.1 Harmonic Model
Using the Python code Imber (Imber:2023; Imber:2024) developed in Plummer&Wang2022; Plummer&Wang2023; Plummer2024, we fit each lightcurve with models of 2nd and 3rd order Fourier modes (or harmonic components), with and without damping. Imber is an open-source code available on GitHub111https://github.com/mkplummer/Imber. The code employs Bayesian methods, specifically dynamic nested sampling (Skilling2004; Skilling2006; Higson2019), to find the best fit for each lightcurve using the Python module Dynesty (Speagle2020). Decomposing lightcurves into their harmonic components provides the capability to directly infer parameters including amplitude, period, phase, and damping ratio for each harmonic.
We modify the harmonic model presented in Plummer2024 by adding a damping term to allow for transient atmospheric dynamics as follows,
@@ 3.2 Fitting Scheme
where denotes inferred parameters, represents observed data, and the applied model. Per Bayesian standard, is the posterior distribution, is the likelihood function, is the prior distribution, and is Bayesian evidence (e.g., Speagle2020).
BIC is computed in the traditional manner (Kass&Raftery1995),
@@ 3.4 Contribution Function
To map the inferred harmonic modes from wavelength-space to pressure-level space in §4, we use a contribution function. Contribution functions show the emitted radiation at each pressure level in wavelength-space and can be computed (Lothringer2018),
where is pressure, is wavelength, is the Planck function, and is the pressure- and wavelength-dependent optical depth.
For this work, we select a cloudless, , Sonora Bobcat (Marley2021) contribution function (see Figures 3 and 4). The temperature and surface gravity for the model are selected based on findings of Gagne2017 and Vos2023. For similar reasons as McCarthy2025, we make the decision to use a cloudless model to probe the deepest observable pressure levels on SIMP J0136. This approach is supported by the likelihood that planetary-mass objects are composed of an admixture of cloudless and cloudy regions, based on a study of 2MASS 1207 b, an object with similar temperature and surface gravity to SIMP J0136 (Zhang2025). It should be noted that we do not necessary expect harmonic parameters to increase or decrease in the same manner as the pressure they probe (i.e., the deepest pressure should not necessarily have to have the shortest periods and lowest amplitudes).
As adapting a single contribution function may bias results (e.g., Figure 4 in Dobbs-Dixon2017), we compared the vertical variation maps described in §4.2.1 to maps created using the retrieved contribution function in Figure 3 of Nasedkin2025. We found that both maps contained the same key features with only minor differences that did not detract from our conclusions.